A long distance runner starts at the beginning of a trail and runs at a rate of
4 miles per hour. Three hours later, a cyclist starts at the beginning of the trail and travels at a rate of
19 miles per hour. What is the amount of time that the cyclist travels before overtaking the runner?
Do not do any rounding.

Answer by richwmiller(17219)

I walk 4 mph!!
Gaining speed 19 -4= 15 mph
Headstart 4*3 =12 miles
Time to catch up 12/15=0.8 hours
Time to catch up =48 minutes
Algebraic solution
4(t+3)=19t
4t+12=19t
12=19t-4t
12=15t
t=0.8 hours
Distance to meeting 0.8*19=15.2 miles